OpenStudy (anonymous):

Evaluate 30 ÷ (6 − 1) • 2 3 8 10 12

OpenStudy (anonymous):

PeMDAS...

OpenStudy (anonymous):

8

OpenStudy (anonymous):

(6-1) = 5 5 x 2 = 10 10/30=3

OpenStudy (anonymous):

30/5 x2 = left to right 6 x 2= 12